OpenStudy (anonymous):

You go to the Whome Insurance Company for auto insurance. They charge $525.00 as a base rate for six months of insurance for your vehicle. They offer a 15% discount for safe driver habits, and a 10% discount for an excellent credit rating or 5% for a good rating. They increase the rate by 5% and 10% respectively for a fair or poor credit rating. You have a fair credit rating and qualify for the safe driver discount. How much does the insurance cost you per month?

OpenStudy (anonymous):

525 - 0.15*525 + 0.05*525 = 525*(1 - 0.15 + 0.05) = 525*(0.9) = $472.50

OpenStudy (31356):

$525.00 X 1.05 for a fair credit rating = $551.25 revised premium $551.25 X 0.85 for safe driver habits = $468.56 final premium So the semi-annual premium is $468.56 divided by 6 = $78.09 In the real world the insurer would impose an installment charge to defer your payments on a monthly basis. Medal?

OpenStudy (31356):

Bangkok is wrong